# WavCraft: Audio Editing and Generation with Natural Language Prompts

### v2

@@ 5.3 Objective evaluation
Evaluation on text-to-audio generation. Table 3 shows the objective results of our WavCraft and the compared methods. WavCraft achieves the best FAD and IS score among the three evaluted models on the AudioCaps evaluation dataset. WavCraft also yields the KL score of 1.68, close to the performance of WavJourney (Liu et al., 2023e).
@@ 5.4 Case study
Case study 1: Audio scriptwriting. WavCraft took two raw audio materials (i.e., the beginning and end of a fan heading to a soccer match field) and a user instruction as inputs. It first analysed the content of input audio and wrote an audio script conditioned on both the user instruction and audio descriptions. The script was written in the format of python coding and applied to allocate diverse modules, such as target source separation, text-to-audio generation models, and the room simulator, for audio editing. The output of activated modules was mixed together in line with the generated audio script. To the best of our knowledge, WavCraft is the only audio agent capable of such complex editing task without an explicit user command.
Case study 2: Human-computer co-creation. We illustrate how WavCraft interacts with an user during the process of audio production. WavCraft started with a basic replacement editing task: replacing the female speech in the audio with another female saying. After the user went through the generated recording, namely OUTPUT1_WAV, the user further instructed the system to drop the audio content between 6-10s. WavCraft kept track of the current chat and generated OUTPUT2_WAV based on the new user instruction and the previous executation lines. Likewise, WavCraft continued to generate OUTPUT3_WAV in response to user’s request about“Add more cheers sound in the end” while taking into consideration the previous instructions. Throughout the process of audio production, in addition to the generated audio, WavCraft also provides users with the executable code together with their comments. We hope this will improve the explainability of WavCraft operations.
@@ 6 Limitations
Despite WavCraft demonstrating a desirable ability of audio editing and generation, there still exists some limitations: 1) Audio analysis: While WavCraft involves an audio analysis module to describe the raw materials, the performance of existing audio analysis models is limited, hindering WavCraft from precisely describing input audio with temporal relationship. 2) Inference speed: WavCraft needs to call diverse APIs to solve a complex task, which introduces time costs during inference. Reducing the inference time will facilitate a more seamless human-AI co-creation, meeting the requirement of more practical applications.

### v3

@@ 5.3 Objective evaluation
Evaluation on text-to-audio generation. Table 4 shows the objective results of our WavCraft and the compared methods. WavCraft achieves the best FAD and IS score among the three evaluted models on the AudioCaps evaluation dataset. WavCraft also yields the KL score of 1.68, close to the performance of WavJourney (Liu et al., 2023d).
5.4 Subjective evaluation
Figure 2: Overall subjective evaluation on audio editing quality by comparing the performance of the ground truth, AUDIT, and the proposed WavCraft.
Figure 2 shows the overall subjective evaluation results by comparing the performance of the ground truth, AUDIT, and our WavCraft. MOS value of WavCraft is very close to that of the ground truth while outperform the AUDIT’s by a large margin.
Figure 4 compares the performance of the ground truth, AUDIT, and the proposed WavCraft in terms of frequency, time, and volume. WavCraft achieves the best results on frequency and time control compared to the AUDIT and even the ground truth. We hypothesise this is partly because the raw audio material used by the ground truth is less perceptually significant than the audio segments generated by WavCraft. In addition, WavCraft achieved a better MOS value compared to AUDIT in terms of volume.
Figure 4 shows the performance of WavCraft and AUDIT$+$ in terms of audio-text relevance, audio coherence, naturalness, engagement, and creativity. The WavCraft yielded a better scores than AUDIT$+$ from all aspects.
Figure 3: Subjective evaluation on the quality of edited audio obtained from the ground truth, AUDIT, and the proposed WavCraft in terms of frequency, temporal, and volume scale.
Figure 4: Comparing the ability of audio storytelling between AUDIT$+$ and the proposed WavCraft in terms of audio-text relevance, audio coherence, naturalness, engagement, and creativity.
@@ 5.5 Case study
Case study 1: Audio scriptwriting. WavCraft takes two raw audio materials (i.e., the beginning and end of an audio titled ”a fan heading to a soccer match field”) and a user instruction as inputs. It first analyses the content of input audio and write an audio script conditioned on both the user instruction and audio descriptions. The script is written in the format of python coding and applied to allocate diverse modules, such as target source separation, text-to-audio generation models, and the room simulator, for audio editing. The output of activated modules is mixed together in line with the generated audio script. To the best of our knowledge, WavCraft is the only audio agent capable of such complex editing task without an explicit user command.
Case study 2: Human-AI co-creation. We illustrate how WavCraft interacts with an user during the process of audio production. WavCraft starts with a basic replacement editing task: replacing the female speech in the audio with another female speaking. After the user went through the generated recording, namely OUTPUT1_WAV, the user further instructed the system to remove the audio content between 6-10s. WavCraft keeps track of the current conversation and generate OUTPUT2_WAV based on the new user instruction and the previous executation lines. Likewise, WavCraft continued to generate OUTPUT3_WAV in response to user’s request about “Add more cheers sound in the end” while taking into consideration the previous instructions. Throughout the process of audio production, in addition to the generated audio, WavCraft also provide users with the executable code together with their comments. We hope this could improve the explainability of WavCraft operations.
@@ 6 Limitations
Despite WavCraft demonstrating a desirable ability of audio editing and generation, there still exists some limitations: 1) Audio analysis: While WavCraft involves an audio analysis module to describe the raw materials, the performance of existing audio analysis models is limited, hindering WavCraft from precisely describing input audio with temporal relationship. 2) Inference cost: WavCraft needs to call diverse APIs to solve a complex task, which introduces time costs during inference. Reducing the inference time will facilitate a more seamless human-AI co-creation, meeting the requirement of more practical applications.
@@ Acknowledgements
Appendix A Prompt template.
Table 5 and Table 6 list the instruction prepended to the user query for the first and following round dialogs, respectively. We exploits the LLM’s ability of in-context learning to allocate diverse audio modules by mimicing previous examples. In the follow-up instruction, we guide LLM to pay attention to the current instruction together with the those of previous dialogs. Therefore, WavCraft can perform audio manipulation by following current request while keeping track of the previous instructions.
Table 5: WavCraft’s predifined instruction prepended to the user query.
INPUT INSTRUCTION
You are an professional audio editor. Try to follow the instruction I give using several predefined tools:
LEN(wav) # returns the duration of ‘wav‘ in seconds
MIX(wavs: list[tuple]) # returns the mixture of the input ‘wavs‘
CAT(wavs: list) # returns the concatenated wav using input ‘wavs‘
SPLIT(wav, break_points=list[float]) # returns the split wavs using ‘break_points‘
ADJUST_VOL(wav, volume: int) # returns the adjusted wav by ‘volume‘
TTA(text: str, length: float, volume: int) # returns a generated audio conditioned on ‘text‘
TTM(text: str, melody, length: float, volume: int) # returns a generated music conditioned on ‘text‘
and (optional) ‘melody‘
TTS(text: str, volume: int) # returns a generated speech conditioned on ‘text‘ and ‘speaker‘. ‘speaker‘
should be in [’Male1_En’, ’Male2_En’, ’Female1_En’, ’Female2_En’]
SR(wav, ddim_steps: int, guidance_scale: float, seed: int) # Returns a wav upsampled to 48kHz
TSS(wav, text: str) # returns foreground and background wav conditioned on ‘text‘
ADD_NOISE(wav, min_snr_db: float, max_snr_db: float)
# returns a generated audio mixed with gaussian noise
LOW_PASS(wav, min_cutoff_freq: float, max_cutoff_freq: float, min_rolloff: int, max_rolloff: int)
# returns a generated audio processed by low pass filter
HIGH_PASS(wav, min_cutoff_freq: float, max_cutoff_freq: float, min_rolloff: int, max_rolloff: int)
# returns a generated audio processed by high pass filter
ADD_RIR(wav, ir) # returns a generated audio mixed with a given room impulse response
ROOM_SIMULATE(wav, min_size_x: float, max_size_x: float, min_size_y: float, max_size_y: float,
min_size_z: float, max_size_z: float, min_absorption_value: float, max_absorption_value: float,
min_source_x: float, max_source_x: float, min_source_y: float, max_source_y: float, min_source_z: float,
max_source_z: float, min_mic_distance: float, max_mic_distance: float, min_mic_azimuth: float,
max_mic_azimuth: float, min_mic_elevation: float, max_mic_elevation: float) # returns a synthesized
audio by mixing the input ‘wav‘ with a room-specific synthesized impulse response
INPAINT(wav, text: str, onset: float, offset: float, duration: float)
# returns a fixed audio where the part between ‘onset‘ and ‘offset‘ has been inpainted
I will give you several examples:
Instruction:
Increase the volume of child speech by 5 dB, decrease the volume of drum by 3 dB, drop the sound of
machine sound.
Code:
# Separate the sound of ’child speech’ from the mixture and return both ’child speech’ and the
background sounds
WAV0, WAV1 = TSS(INPUT_WAV0, text=”child speech”)
# Separate the sound of ’drum’ from the mixture and return both ’drum’ and the background sounds
WAV2, WAV3 = TSS(WAV1, text=”drum”)
# Drop the sound of ’machine sound’ from the mixture
_, WAV3 = TSS(WAV3, text=”machine sound”)
# Increace the volume of ”child speech” by 5dB
WAV0 = ADJUST_VOL(WAV0, volume=5)
# Decrease the volume of ’drum’ by 5dB
WAV2 = ADJUST_VOL(WAV2, volume=-3)
# Mix the resulted sounds together
OUTPUT_WAV = MIX([(WAV0, 0), (WAV2, 0), (WAV3, 0)])
INPUT INSTRUCTION (con’d)
Instruction:
Extract 1-5s of the first audio with a low-pass filter to simulate the sound coming from inside a building.
Replace male speech with dog barking in the second audio. Upsample the mix.
Code:
# Truncate the sound between 1s and 5 s
_, WAV0, _ = SPLIT(INPUT_WAV0, break_points=[1, 5])
# Add a low-pass filter
WAV0 = LOW_PASS(WAV0, min_cutoff_freq=300.0, max_cutoff_freq=800.0,
min_rolloff=6, max_rolloff=12)
# Extract the sound of ’male speech’ from the truncated sound
WAV1, WAV2 = TSS(INPUT_WAV1, text=”male speech”)
# Generate the sound of ’dog barking’ with the same length with the sound of ’male speech’
WAV3 = TTA(text=”dog barking”, length=LEN(WAV1), volume=4)
# Combine the sounds by mixing them together
MIXTURE_WAV = MIX([(WAV3, 0), (WAV2, 0), (WAV0, 0)])
# Perform super-resolution on the mixture of sounds
OUTPUT_WAV = SR(MIXTURE_WAV)
Instruction:
Isolate train sound in the input audio, apply a high-pass filter and increase the volume by 3 dB.
Repeat it five times to simulate a longer train passing.
Code:
# Extract the sound of a train from the audio
WAV0, _ = TSS(INPUT_WAV0, text=”train”)
# Apply a high-pass filter to reduce low-frequency noise
FILTERED_WAV0 = HIGH_PASS(WAV0, min_cutoff_freq=500.0, max_cutoff_freq=1000.0,
min_rolloff=6, max_rolloff=12)
# Increase the volume by 3 dB
FILTERED_WAV0 = ADJUST_VOL(FILTERED_WAV0, volume=3)
# Concatenate the filtered train sound three times
OUTPUT_WAV = CAT([FILTERED_WAV0] * 5)
Instruction:
Extract the hammer sound from the first audio, and truncate it from the start towards 2 second.
Remove the sound of baby crying in the second audio, and then decrease the volume by 1 dB.
Mix two audio together, and the second sound should begin from 1 second. Add a reverb effect
to the mixture sound using the third audio.
Code:
# Extract the hammer sound from the first audio
WAV0, _ = TSS(INPUT_WAV0, text=”hammer”)
# Truncate from the start towards 2 second
WAV0, _ = SPLIT(WAV0, break_points=[2])
# Drop the sound of baby crying in the second audio
_, WAV1 = TSS(INPUT_WAV1, text=”baby crying”)
# Decrease the volume by 1 dB
WAV1 = ADJUST_VOL(WAV1, volume=-1)
# Mix the ouput sounds together
MIXED_WAV = MIX([(WAV0, 0), (WAV1, 1)])
# Add a reverb effect using room impulse response
OUTPUT_WAV = ADD_RIR(MIXED_WAV, ir=INPUT_WAV2)
INPUT INSTRUCTION (con’d)
Instruction:
Inpaint the first audio between 2s and 5s with the text ”a car passing by with rain falling”.
Generate a 10s long jazz music piece with the second audio as melody, then mix it with the
sound of rain from the first, starting at 3s into the jazz music.
Code:
# Inpaint the first audio between 2s and 5s with the text ”a car passing by with rain falling”
WAV0 = INPAINT(INPUT_WAV0, text=”a car passing by with rain falling”, onset=2,
offset=5, duration=LEN(INPUT_WAV0))
# Generate a 10-second jazz music piece
WAV1 = TTM(text=”jazz”, melody=INPUT_WAV1, length=10.0, volume=5)
# Extract the sound of rain from the audio file
WAV0, _ = TSS(WAV0, text=”rain”)
# Mix the jazz music with the rain sound, starting the rain at 3 seconds
OUTPUT_WAV = MIX([(WAV0, 0), (WAV1, 3)])
Instruction:
Remove wind sound from an outdoor recording. Generate a 5-second saxophone music
with happy mood followed by ”Bravo”. Mix the generated sound with the outdoor
recording and simulate the mixture in a small room with high absorption.
Code:
# Drop the sound of wind from the original recording
_, WAV0 = TSS(INPUT_WAV0, text=”wind”)
# Generate a 5-second saxophone music with happy mood followed by a male speech
”Bravo”.
WAV1 = TTM(text=”happy saxophone”, length=5.0, volume=4)
# Generate a speech ”Bravo”
WAV2 = TTS(”Bravo”, volume=5)
# Concatenate the generated sound together
CONCAT_WAV = CAT([WAV1, WAV2])
# Mix the generated sound with the background sound
MIXED_WAV = MIX((WAV0, 0), (CONCAT_WAV, 0))
# Simulate the recording in a small room with high absorption
OUTPUT_WAV = ROOM_SIMULATE(MIXED_WAV, min_size_x=3, max_size_x=4,
min_size_y=3, max_size_y=4, min_size_z=2.5, max_size_z=3, min_absorption_value=0.7,
max_absorption_value=0.9, min_source_x=1, max_source_x=1.5, min_source_y=1,
max_source_y=1.5, min_source_z=1, max_source_z=1.5, min_mic_distance=1,
max_mic_distance=1.5, min_mic_azimuth=45, max_mic_azimuth=90,
min_mic_elevation=20, max_mic_elevation=30)
Table 6: WavCraft’s follow-up instruction to make sure the consistency of generated audio within the multi-round dialog between human and AI.
FOLLOW-UP INSTRUCTION
Regenerate the code by appending the new instruction to the previous instructions.
The code must start with the provided audio (e.g., INPUT_WAV0) and cannot take
the output from previous phase (i.e., ‘OUTPUT_WAV‘) as a known input. The new
instruction is: